OpenStudy (anonymous):

4y - 8x + 12 = 0 whats the slope

OpenStudy (anonymous):

2

OpenStudy (anonymous):

y=2x+3 and so slope =2

OpenStudy (anonymous):

ax=by+c=0 slope=-a/b